What is OH- of a solution whose pH is 4.732

1
Expert's answer
2022-05-24T06:00:03-0400

Solution:

For any aqueous solution at 25C:

pH + pOH = 14

pOH = 14 − pH = 14 − 4.732 = 9.268

pOH = 9.268


We can convert between pOH and [OH] using the following equation:

pOH = −log[OH]

[OH] = 10−pOH = 10−9.268 = 5.395×10−10

[OH] = 5.395×10−10 M


Answer: The hydroxide ion (OH) concentration is 5.395×10−10 M